# Can repeating and non-repeating FRBs be drawn from the same population?

### v1

@@ 6 Conclusions
The applicability of Zipf’s law to FRBs is derived directly from observational data. It implies that the effective distribution of intrinsic repetition rates (above a fixed energy) follows a PL: $P(>\mathcal{R}_{*})\propto\mathcal{R}_{*}^{-a}$ with $a$ close to unity. We explored the implications of such a distribution for the observable properties of FRBs detected in wide-field surveys such as ASKAP, DSA, and CHIME. Specifically, we computed the expected distributions of fluence, distance, and observed repetition number as functions of $a$ and $\gamma$ (where $\gamma$ is the PL index describing the burst energy distribution from individual sources). Our analysis shows that the behavior of the FRB population can be qualitatively divided into four distinct regimes in the $a-k$ parameter space (with $k$ denoting the number of observed bursts from a source), depending on the relationship between $a,k$, and the critical value $a_{\gamma}=k_{\gamma}=3/(2(\gamma-1))$. For typical observed values - $\gamma\approx 1.7$ and $a\approx 1$ repeaters are dominated by a small number of highly active sources, while non-repeaters include comparable contributions from both rare, active sources and more common, intrinsically inactive ones. All observed FRBs tend to originate from distant sources, although repeaters are, on average, slightly closer than non-repeaters (consistent with trends in dispersion measures; see Fig. 5). A much sharper distinction exists between sources that are observed at least once and those that are never detected. The latter are predicted to be far more numerous and significantly less active. For example, the corrected repetition rate of SGR 1935+2154 (accounting for the observed energy of its bursts) yields an expected detection rate (above CHIME threshold) of only $\sim few\times 10^{-9}$ bursts per such source at cosmological distances—comparable to the inferred ratio of CHIME-detected sources to the estimated number of inactive, SGR-like sources in the Universe. This apparent coincidence emerges naturally if the source repetition rate distribution follows a power law with index $a\approx 1.1-1.3$. Notably, the same distribution also reproduces the observed ratio between CHIME repeaters and apparent non-repeaters. Since $a-1\ll 1$, the relative fraction of repeaters increases only very slowly with increased observation time or improved sensitivity (as $N_{\geq 2}/N_{\geq 1}\propto(\Phi_{\rm lim}^{1-\gamma}T)^{a-1}$ see §A and Fig. 5). This weak dependence could be misinterpreted as evidence that repeaters and non-repeaters represent fundamentally different source populations, even if all FRBs originate from the same underlying class.
@@
A key strength of our analysis is its independence from any specific progenitor or emission model. Nonetheless, the consistency of the least active FRB sources with known Galactic SGRs, combined with the success of a single PL of the activity-rate distribution in explaining the full observed population, naturally points to magnetars as the dominant—or possibly sole—origin of FRBs. What then accounts for the wide variation in inferred source densities and activity levels among different magnetars? Beniamini & Kumar (2023, 2025) suggest that the inclination between spin and magnetic axes ($\alpha$) can play a major role in understanding various apparent differences between repeaters and non-repeaters (including rate, source densities and energetics as relevant for the present analysis, but also resolving the lack of rotational periodicity in arrival time data of active repeaters, explaining the lack of polarization angle in active repeaters compared to polarization angle swing seen in some non-repeaters, providing a geometrical explanation for the CHIME spectro-temporal dichotomy and explaining the periodic modulation within the duration of the long non-repeating FRB 20191221A). This raises the question - what fraction of the spread in inferred source densities can be accounted for by magnetic inclination? The fraction of the sky that is illuminated by a given misaligned NS’s polar cap is $\propto\sin\alpha\approx\alpha$. The intrinsic probability of the inclination being $\leq\alpha$, $P_{\rm int}(<\alpha)$, depends on the initial distribution and on potential alignment that may take place as the magnetar evolves, both of which are uncertain. Therefore, all else being equal, the probability to observe a NS with inclination smaller than $P_{\rm ob}(<\alpha)$ is $\sim P_{\rm int}(<\alpha)\alpha$. An illustrative example is the case of a purely isotropic distribution. In this case we have $P_{\rm int}\propto\alpha^{2}$ and $P_{\rm ob}(<\alpha)\propto\alpha^{3}$. This holds for $\rho\lesssim\alpha\lesssim\pi/2$ where $\rho$ is the half opening angle of the polar cap in the observer frame. Considering likely values of the spin period, one might expect $10^{-2.5}\lesssim\rho\lesssim 10^{-1.5}$. Therefore, it is plausible that $\alpha^{3}$ varies by up to $\sim 7.5$ orders of magnitude and that inclination plays a significant and perhaps even the major rule in explaining the inferred source density variations. The rotation of the emission beam also affects the observed rate, which will increase with decreasing $\alpha$ according to $\mathcal{R}_{\rm ob}\propto\alpha^{-1}\mathcal{R}_{\rm int}(\alpha)$, where $\mathcal{R}_{\rm int}(\alpha)$ is the intrinsic dependence of the rate on $\alpha$. Agreement with the inferred $a\approx 1.1-1.3$ requires (for an isotropic inclination distribution) that $\mathcal{R}_{\rm int}(\alpha)\mathrel{\raisebox{3.01385pt}{\mbox{$\underset{\vbox{\hbox{$\sim$}}}{\vbox{\hbox{$\propto$}}}$}}}\alpha^{-1.8}-\alpha^{-1.3}$. Whether or not this can be reconciled depends on uncertain bursting properties and is an interesting topic for future investigations. Variation in other properties such as magnetic field strength or age—as well as temporal variations in burst rate, energy distribution slope $\gamma$, or maximum burst energy—can also partially contribute to the intrinsic rate - source density Zipf law. Finally, an alternative possibility relates to the presence of a rare sub-population of ultra-long-period magnetars that might be especially prolific FRB emitters (see e.g. Beniamini et al. 2020, 2023). Note that in this case, the polar cap angle can be much smaller than mentioned above, making it easier to explain a large range in rates and source densities, even without a strong dependence of $P_{\rm int}(<\alpha),\mathcal{R}_{\rm int}(\alpha)$ on $\alpha$.
@@ ACKNOWLEDGEMENTS
We thank Zorawar Wadiasingh, Om Gupta and Rob Fender for helpful comments on the manuscript. The work was funded in part by an NSF grant AST-2009619 (PK), a NASA grant 80NSSC24K0770 (PK and PB), a grant (no. 2020747) from the United States-Israel Binational Science Foundation (BSF), Jerusalem, Israel (PB) and by a grant (no. 1649/23) from the Israel Science Foundation (PB).
@@ Appendix A Numerical verification
where $A_{\rm src}^{\rm sfr}$ is the conversion fraction from star formation rate to population density of FRB sources. If the latter are magnetars, then approximately, $A_{\rm src}^{\rm sfr}\approx 20\mbox{yr}M_{\odot}^{-1}$ (corresponding to roughly 30 magnetars in the Milky Way and a number density of Milky-Way like galaxies of $\sim 0.01\mbox{ Mpc}^{-3}$). Some recent studies (Qiang et al., 2022; Zhang & Zhang, 2022; Lin & Zou, 2024; Gupta et al., 2025; Horowicz & Margalit, 2025; Acharya & Beniamini, 2025) have suggested that FRBs arising from a mix of SFR and stellar mass following populations, fit better with the observed redshift and host galaxy distributions. Such a possibility is straightforward to include in this analysis. Since its effect on the results of this study are minor and in order to avoid introducing additional weakly constrained parameters, we ignore this complication in the estimates below. We assume also that the average spectrum of FRBs can be characterized using $E_{\nu}\propto\nu^{\alpha}$. Previous studies (Macquart et al., 2019; Shin et al., 2023) have found $\alpha\approx-1.5$. The measured fluence of an FRB is then given by $\Phi=E_{\nu_{0}}(1+z)^{\alpha}/(4\pi r^{2})$ where $\nu_{0}$ is the frequency at which the fluence is measured, $r$ is the proper distance and the factor $(1+z)^{\alpha}$ is a $k$-correction, i.e. converting from the measured $\nu_{0}$ to the corresponding frequency in the comoving frame $\nu=\nu_{0}(1+z)$.
@@
where $\Omega_{\rm srv}/4\pi$ is the fraction of the sky viewed by the survey (approximately 0.5 for CHIME), $\Phi_{\rm lim}$ is the limiting (spectral) fluence of the survey and $E_{\rm lim}(z)=4\pi r^{2}\Phi_{\rm lim}(1+z)^{-\alpha}$. $d\mbox{P}/d\mathcal{R}_{\rm*}\propto\mathcal{R}_{\rm*}^{-a-1}$ is the probability of a source having a given intrinsic rate (at $E_{\rm*}$). Pois$(k|\lambda)$ is the Poisson probability distribution of observing $k$ events considering an expected value of $\lambda$. The value $\lambda=\frac{T}{1+z}\mathcal{R}_{\rm*}\left[\frac{E_{\rm lim}(z)}{E_{\rm*}}\right]^{1-\gamma}$ accounts for the rate of the source above the minimum energy detectable by the survey and considering the (comoving) time that it has been observed by the survey.
@@
The results of numerical integration for $N_{\geq 2}/N_{\geq 1},N_{\geq 1}/N_{\geq 0}$ as a function of $a$ were shown in Fig. 3 where it was shown that $1.1\lesssim a\lesssim 1.3$ can simultaneously match the observed $N_{\geq 1}/N_{\geq 0}$, $N_{\geq 2}/N_{\geq 1}$ and $\lambda_{\rm 0,lim}$. Fig. 4 shows how the ratio of repeaters to total detected sources, $N_{\geq 2}/N_{\geq 1}$, evolves for a fixed value of $a=1.1$ , as a function of decreasing the fluence threshold or increasing the average observation time per sky location. For a small value of $a-1$, the analytic expression $N_{\geq 2}/N_{\geq 1}\propto(\Phi_{\rm lim}^{1-\gamma}T)^{a-1}$ (Eq. 29) predicts that the repeater fraction depends only weakly on both limiting fluence and observing time. This weak dependence is confirmed by our numerical results: increasing the observation time by a factor of 10 leads to only a modest increase - by a factor of $\sim 1.8$ in the fraction of repeaters. Reducing the fluence threshold by a similar factor yields an even smaller effect. The figure also illustrates how the repeater fraction varies across the sky, based on the relation between exposure time and declination in the first CHIME catalog. Since CHIME observes higher declinations for longer, the fraction of repeaters is slightly elevated in those regions. However, this variation is relatively small, again reflecting the weak dependence of the repeater fraction on exposure (as a rough approximation, CHIME’s exposure distribution roughly satisfies $T\propto(\pi/2-\delta)^{-1}$, CHIME/FRB Collaboration et al. 2021, leading to $N_{\geq 2}/N_{\geq 1}\propto(\pi/2-\delta)^{1-a}$). We also show, in Fig. 5. the distribution of distances, external dispersion measures and intrinsic rates for both observed repeaters and non-repeaters. As shown in Fig. 2, for $a=1.1$, both repeaters and non-repeaters are predominantly detected at cosmological distances. Interestingly, repeaters are, on average, a few tens of percent closer than non-repeaters. This leads to typical intergalactic dispersion measures of $\mbox{DM}_{\rm ext,rep}\sim 400\mbox{pc cm}^{-3}$ for repeaters, compared to $\mbox{DM}_{\rm ext,NR}\sim 670\mbox{pc cm}^{-3}$ for non-repeaters—both consistent with observed values (Chime/Frb Collaboration et al., 2023). This agreement is particularly striking given that our modeling was developed independently of dispersion measure data, using a simple framework and approximate detectability criteria. While the two populations show some overlap in distance, they differ much more clearly in their intrinsic repetition rates: repeaters are dominated by the rarest and most active sources, whereas the non-repeater population includes a roughly equal mix of common, inactive sources and rare, highly active ones.
@@
Appendix B Correlation between rate and maximal energy

### v2

@@ 6 Conclusions
The applicability of Zipf’s law to FRBs is derived directly from observational data. It implies that the effective distribution of intrinsic repetition rates (above a fixed energy) follows a PL: $P(>\mathcal{R}_{*})\propto\mathcal{R}_{*}^{-a}$ with $a$ close to unity. We explored the implications of such a distribution for the observable properties of FRBs detected in wide-field surveys such as ASKAP, DSA, and CHIME. Specifically, we computed the expected distributions of fluence, distance, and observed repetition number as functions of $a$ and $\gamma$ (where $\gamma$ is the PL index describing the burst energy distribution from individual sources). Our analysis shows that the behavior of the FRB population can be qualitatively divided into four distinct regimes in the aka-k parameter space (with $k$ denoting the number of observed bursts from a source), depending on the relationship between $a,k$, and the critical value $a_{\gamma}=k_{\gamma}=3/(2(\gamma-1))$. For typical observed values - $\gamma\approx 1.7$ and $a\approx 1$ - repeaters are dominated by a small number of highly active sources, while non-repeaters include comparable contributions from both rare, active sources and more common, intrinsically inactive ones. All observed FRBs tend to originate from distant sources, although repeaters are, on average, slightly closer than non-repeaters (consistent with trends in dispersion measures; see Fig. 5). A much sharper distinction exists between sources that are observed at least once and those that are never detected. The latter are predicted to be far more numerous and significantly less active. For example, the corrected repetition rate of SGR 1935+2154 (accounting for the observed energy of its bursts) yields an expected detection rate (above CHIME threshold) of only $\sim few\times 10^{-9}$ bursts per such source at cosmological distances - comparable to the inferred ratio of CHIME-detected sources to the estimated number of inactive, SGR-like sources in the Universe. This apparent coincidence emerges naturally if the source repetition rate distribution follows a power law with index $a\approx 1.1-1.3$. Notably, the same distribution also reproduces the observed ratio between CHIME repeaters and apparent non-repeaters. Since $a-1\ll 1$, the relative fraction of repeaters increases only very slowly with increased observation time or improved sensitivity (as $N_{\geq 2}/N_{\geq 1}\propto(\Phi_{\rm lim}^{1-\gamma}T)^{a-1}$ see §A and Fig. 5). This weak dependence could be misinterpreted as evidence that repeaters and non-repeaters represent fundamentally different source populations, even if all FRBs originate from the same underlying class.
@@
A key strength of our analysis is its independence from any specific progenitor or emission model. Nonetheless, the consistency of the least active FRB sources with known Galactic SGRs, combined with the success of a single PL of the activity-rate distribution in explaining the full observed population, naturally points to magnetars as the dominant—or possibly sole—origin of FRBs. What then accounts for the wide variation in inferred source densities and activity levels among different magnetars? Beniamini & Kumar (2023, 2025) suggest that the inclination between spin and magnetic axes ($\alpha$) can play a major role in understanding various apparent differences between repeaters and non-repeaters (including rate, source densities and energetics as relevant for the present analysis, but also resolving the lack of rotational periodicity in arrival time data of active repeaters, explaining the lack of polarization angle swing in active repeaters compared to polarization angle swing seen in some non-repeaters and providing a geometrical explanation for the CHIME spectro-temporal dichotomy). This raises the question - what fraction of the spread in inferred source densities can be accounted for by magnetic inclination? The fraction of the sky that is illuminated by a given misaligned NS’s polar cap is $\propto\sin\alpha\approx\alpha$. The intrinsic probability of the inclination being $\leq\alpha$, $P_{\rm int}(<\alpha)$, depends on the initial distribution and on potential alignment that may take place as the magnetar evolves, both of which are uncertain. Therefore, all else being equal, the probability to observe a NS with inclination smaller than $P_{\rm ob}(<\alpha)$ is $\sim P_{\rm int}(<\alpha)\alpha$. An illustrative example is the case of a purely isotropic distribution. In this case we have $P_{\rm int}\propto\alpha^{2}$ and $P_{\rm ob}(<\alpha)\propto\alpha^{3}$. This holds for $\rho\lesssim\alpha\lesssim\pi/2$ where $\rho$ is the half opening angle of the polar cap in the observer frame. Considering likely values of the spin period, one might expect $10^{-2.5}\lesssim\rho\lesssim 10^{-1.5}$. Therefore, it is plausible that $\alpha^{3}$ varies by up to $\sim 7.5$ orders of magnitude and that inclination plays a significant and perhaps even the major rule in explaining the inferred source density variations. The rotation of the emission beam also affects the observed rate, which will increase with decreasing $\alpha$ according to $\mathcal{R}_{\rm ob}\propto\alpha^{-1}\mathcal{R}_{\rm int}(\alpha)$, where $\mathcal{R}_{\rm int}(\alpha)$ is the intrinsic dependence of the rate on $\alpha$. Agreement with the inferred $a\approx 1.1-1.3$ requires (for an isotropic inclination distribution) that $\mathcal{R}_{\rm int}(\alpha)\mathrel{\raisebox{3.01385pt}{\mbox{$\underset{\vbox{\hbox{$\sim$}}}{\vbox{\hbox{$\propto$}}}$}}}\alpha^{-1.8}-\alpha^{-1.3}$. Whether or not this can be reconciled depends on uncertain bursting properties and is an interesting topic for future investigations. Variation in other properties such as magnetic field strength or age-as well as temporal variations in burst rate, energy distribution slope $\gamma$, or maximum burst energy-can also partially contribute to the intrinsic rate - source density Zipf law. Finally, an alternative possibility relates to the presence of a rare sub-population of ultra-long-period magnetars that might be especially prolific FRB emitters (see e.g. Beniamini et al. 2020, 2023). Note that in this case, the polar cap angle can be much smaller than mentioned above, making it easier to explain a large range in rates and source densities, even without a strong dependence of $P_{\rm int}(<\alpha),\mathcal{R}_{\rm int}(\alpha)$ on $\alpha$.
@@ ACKNOWLEDGEMENTS
We thank Zorawar Wadiasingh, Om Gupta, Rob Fender and the anonymous referee for helpful comments on the manuscript. The work was funded in part by an NSF grant AST-2009619 (PK), a NASA grant 80NSSC24K0770 (PK and PB), a grant (no. 2020747) from the United States-Israel Binational Science Foundation (BSF), Jerusalem, Israel (PB) and by a grant (no. 1649/23) from the Israel Science Foundation (PB).
[…]
@@ Appendix A Numerical verification
where $A_{\rm src}^{\rm sfr}$ is the conversion fraction from star formation rate to population density of FRB sources. If the latter are magnetars, then approximately, $A_{\rm src}^{\rm sfr}\approx 20\mbox{yr}M_{\odot}^{-1}$ (corresponding to roughly 30 magnetars in the Milky Way and a number density of Milky-Way like galaxies of $\sim 0.01\mbox{ Mpc}^{-3}$). Some recent studies (Qiang et al., 2022; Zhang & Zhang, 2022; Lin & Zou, 2024; Gupta et al., 2025; Horowicz & Margalit, 2025; Acharya & Beniamini, 2025) have suggested that FRBs arising from a mix of SFR and stellar mass following populations, fit better with the observed redshift and host galaxy distributions. Such a possibility is straightforward to include in this analysis. Since its effect on the results of this study are minor and in order to avoid introducing additional weakly constrained parameters, we ignore this complication in the estimates below. We assume also that the average spectrum of FRBs can be characterized using $E_{\nu}\propto\nu^{\alpha}$. The measured fluence of an FRB is then given by $\Phi=E_{\nu_{0}}(1+z)^{\alpha}/(4\pi r^{2})$ where $\nu_{0}$ is the frequency at which the fluence is measured, $r$ is the proper distance and the factor $(1+z)^{\alpha}$ is a $k$-correction, i.e. converting from the measured $\nu_{0}$ to the corresponding frequency in the comoving frame $\nu=\nu_{0}(1+z)$.
@@
$N_{\geq k}=\frac{\Omega_{\rm srv}}{4\pi}\int_{0}^{\infty}dT\int_{\mathcal{R}_{\rm 0,*}}^{\infty}d\mathcal{R}_{\rm*}\int_{0}^{\infty}dV(z)\frac{d\mbox{P}}{dT}\frac{d\mbox{P}}{d\mathcal{R}_{\rm*}}n_{\rm src}(z)\mbox{Pois}\left(k\bigg{|}\frac{T}{1+z}\mathcal{R}_{\rm*}\left[\frac{E_{\rm lim}(z)}{E_{\rm*}}\right]^{1-\gamma}\right)\Theta(E_{\rm c}-E_{\rm lim}(z))$ (A2)
@@
The results of numerical integration for $N_{\geq 2}/N_{\geq 1},N_{\geq 1}/N_{\geq 0}$ as a function of $a$ were shown in Fig. 3 where it was shown that $1.1\lesssim a\lesssim 1.3$ can simultaneously match the observed $N_{\geq 1}/N_{\geq 0}$, $N_{\geq 2}/N_{\geq 1}$ and $\lambda_{\rm 0,lim}$. Fig. 4 shows how the ratio of repeaters to total detected sources, $N_{\geq 2}/N_{\geq 1}$, evolves for a fixed value of $a=1.1$ , as a function of decreasing the fluence threshold or increasing the average observation time per sky location. For a small value of a1a-1, the analytic expression $N_{\geq 2}/N_{\geq 1}\propto(\Phi_{\rm lim}^{1-\gamma}T)^{a-1}$ (Eq. 29) predicts that the repeater fraction depends only weakly on both limiting fluence and observing time. This weak dependence is confirmed by our numerical results: increasing the observation time by a factor of 10 leads to only a modest increase - by a factor of $\sim 1.8$ - in the fraction of repeaters. Reducing the fluence threshold by a similar factor yields an even smaller effect. The figure also illustrates how the repeater fraction varies across the sky, based on the relation between exposure time and declination in the first CHIME catalog. Since CHIME observes higher declinations for longer, the fraction of repeaters is slightly elevated in those regions. However, this variation is relatively small, again reflecting the weak dependence of the repeater fraction on exposure (as a rough approximation, CHIME’s exposure distribution roughly satisfies $T\propto(\pi/2-\delta)^{-1}$, CHIME/FRB Collaboration et al. 2021, leading to $N_{\geq 2}/N_{\geq 1}\propto(\pi/2-\delta)^{1-a}$). We also show, in Fig. 5. the distribution of distances, external dispersion measures and intrinsic rates for both observed repeaters and non-repeaters. As shown in Fig. 2, for $a=1.1$, both repeaters and non-repeaters are predominantly detected at cosmological distances. Interestingly, repeaters are, on average, a few tens of percent closer than non-repeaters. This leads to typical intergalactic dispersion measures of $\mbox{DM}_{\rm ext,rep}\sim 400\mbox{pc cm}^{-3}$ for repeaters, compared to $\mbox{DM}_{\rm ext,NR}\sim 670\mbox{pc cm}^{-3}$ for non-repeaters—both consistent with observed values (Chime/Frb Collaboration et al., 2023). This agreement is particularly striking given that our modeling was developed independently of dispersion measure data, using a simple framework and approximate detectability criteria. While the two populations show some overlap in distance, they differ much more clearly in their intrinsic repetition rates: repeaters are dominated by the rarest and most active sources, whereas the non-repeater population includes a roughly equal mix of common, inactive sources and rare, highly active ones.
@@
Appendix B Inferred densities and source activity rates
The nearest source that is a member of a given class of objects and is within the field observable by the telescope, would be the brightest and easiest to detect [^12] Beaming corrections can be important in order to translate the inferred source densities derived in this way to true space densities, but since it is apriori unknown, it is more robust to report the density of sources that are beamed towards the observer. Therefore, if the closest known member of a class of sources is at a distance $d_{\rm p}$ (for cosmological sources, $d_{\rm p}$ here is the proper distance), then the volume within which there is one source is $V_{\rm obs}=\Omega_{\rm srv}d_{\rm p}^{3}/3$ (where $\Omega_{\rm srv}$ is the solid angle covered by telescopes that could have detected such sources, and $\Omega_{\rm srv}\approx 2\pi$ for CHIME). The mean number of sources in the volume is given by $\lambda_{\rm obs}=V_{\rm obs}n_{\rm obs}$ where $n_{\rm obs}$ is the volumetric density of sources. From Poisson statistics, the probability of observing one source in this volume, given $\lambda_{\rm obs}$, is $P(1|\lambda_{\rm obs})=\lambda_{\rm obs}\exp{[-\lambda_{\rm obs}]}$. Using Bayes’ theorem, this can be converted to the probability that the mean number of sources in the volume is $\lambda_{\rm obs}$ given that one source was detected within the volume,
$\frac{dP(\lambda_{\rm obs}|1)}{d\lambda_{\rm obs}}\propto P(1|\lambda_{\rm obs})\frac{dP}{d\lambda_{\rm obs}}.$ (B1)
To progress from here, one needs to make a choice regarding the prior of $dP/d\lambda_{\rm obs}$. We take $dP/d\lambda_{\rm obs}\propto\lambda_{\rm obs}^{-1}$ such that $P(<\lambda_{\rm obs})\propto\log(\lambda_{\rm obs})$ and no apriori bias is introduced in favor of large or small values of $\lambda_{\rm obs}$ (see e.g. Lu et al. 2022 and references therein). Plugging this in Eq. B1, switching variables from $\lambda_{\rm obs}$ to $n_{\rm obs}$ and integrating we get
$P(<n_{\rm obs}|1)=1-\exp[-V_{\rm obs}n_{\rm obs}].$ (B2)
Eq. B2 shows that the median value of $n_{\rm obs}$ is $0.7V_{\rm obs}^{-1}$ and that the 90% confidence intervals are $n_{\rm obs}\in[0.051,3]V_{\rm obs}^{-1}$. We use these values to estimate the volumetric densities of different FRB source types in figure 1.
To calculate the rates of different sources, we have collected data reported in the literature. Since the activity rates of FRB sources may vary significantly over time, we have chosen, wherever possible, the observations corresponding to the longest monitoring of any given source. Errors due to time variation are calculated assuming Poisson statistics. Furthermore, since rates of sources depend on both burst energy and observed frequency, we present all rates in figure 1 at a common energy ($10^{28}\mbox{erg Hz}^{-1}$) and frequency (600MHz - the midpoint of the CHIME band). When observations at those energy and frequency are not available, we take the closest observed energy/frequency and extrapolate using Eq. 1 for the energy and the ‘average’ spectrum $E_{\nu}\propto\nu^{\alpha}$. This ‘averaged spectrum’ takes into account that there is large variation in individual FRB spectra, some being narrow band, some broad band and some spectrally complex, but as shown by e.g. Macquart et al. (2019), when averaged over a larger sample, the fluence is reasonably represented by a power-law behavior. Using the average spectrum and energy dependent rate, if we know the rate of bursts with energy $E_{1}$ at some $\nu_{1}$, $\mathcal{R}(>E_{1}|\nu_{1})$, then the source would on average have the same rate of bursts at $\nu_{0}$ and above $E_{0}=E_{1}(\nu_{0}/\nu_{1})^{\alpha}$ (this is straightforward for wide band bursts - each burst with energy $E_{1}$ at $\nu_{1}$ has an energy $E_{0}$ at $\nu_{0}$. It is therefore the most self-consistent definition of $\alpha$ that guarantees the results do not depend on whether individual bursts are wide or narrow band). With this assumption, we can get the rate of bursts above any other value of $E_{\nu}$ via
$\mathcal{R}(>\!E_{\nu}|\nu_{0})=\mathcal{R}(>\!E_{0}|\nu_{0})\bigg{(}\frac{E_{\nu}}{E_{0}}\bigg{)}^{1-\gamma}=\mathcal{R}(>\!E_{1}|\nu_{1})\bigg{(}\frac{E_{\nu}}{E_{0}}\bigg{)}^{1-\gamma}=\mathcal{R}(>\!E_{1}|\nu_{1})\bigg{(}\frac{E_{\nu}}{E_{1}}\bigg{)}^{1-\gamma}\bigg{(}\frac{\nu_{1}}{\nu_{0}}\bigg{)}^{\alpha(1-\gamma)}$ (B3)
When observations of a given source are not available at 600 MHz and $10^{28}\mbox{erg Hz}^{-1}$, Eq. B3 is used to extrapolate the rates. The errors are extrapolated in the same way, and in addition we allow for uncertainties in the average values of $\gamma$ and $\alpha$. Unless, reported otherwise for a specific source, we take $\gamma=1.7\pm 0.2$ and $\alpha=-1.5\pm 0.3$. Uncertainties in figure 1 mark 90% confidence intervals. We generally choose observations at energies / frequencies as close as possible to our reference values, in order to minimize errors due to the extrapolation.
We include in Fig.1 a calculation of the equivalent volumetric source density and rate that would be associated with CHIME detected sources under the assumption (disfavored in this work) that there is a single activity rate per source underlying all sources (as per §2). These are calculated in the following way. From the number of sources detected twice and those detected once, we can calculate the number of all sources (i.e. including those which have not been detected at all). This can be done in an approximate way using Eq. 14 and correcting for the fraction of sky visible by CHIME (see, Eq. 23), or in a slightly more accurate way using Eq. A2 with $dP/d\mathcal{R}_{*}=\delta(\mathcal{R}_{*}-\mathcal{R}_{0,*})$. Once the total number of sources, $N_{\geq 0}^{\rm sgl}$, has been calculated, we use Eq. A1 to find the corresponding source density at $z=0$. Dividing the all-sky equivalent rate of bursts discovered by CHIME (as reported in the 1st catalog, CHIME/FRB Collaboration et al. 2021) by $N_{\rm\geq 0,CH}^{\rm all-sky}$ we obtain the average rate per source under these assumptions. The latter is then extrapolated from $E_{\rm lim}$ to $10^{28}\mbox{erg Hz}^{-1}$ using Eq. 1 and $\gamma=1.7\pm 0.2$ as described in the previous paragraph. We emphasize that these “effective" densities and rates are derived using a proof by contradiction approach: they assume that all the different sources observed by CHIME can be described by a single characteristic activity rate and typical source density. However, as illustrated by the wide spread of black points in the parameter space in Fig. 1, this assumption is clearly inconsistent with the data based on the calculation of the rates and source densities of individually well monitored sources.
While using the technique outlined at the top of this appendix, $n_{\rm obs}$ is only well estimated by the nearest member of a given observationally defined sub-class of sources (e.g. one with a given repetition rate or other clearly identifiable observable distinctions), the individual source distances ($d_{\rm p}$) maintain their physical meaning also for non-nearest members. An additional advantage of considering the source distances rather than source densities is that, as explained above, in order to estimate $n_{\rm obs}$, it is crucial to have an estimate of the sky coverage of the detecting telescope in the observing mode that detected the burst. This issue is particularly important for non-repeaters detected by instruments other than CHIME, which typically probe only a small fraction of the sky (leading to a large and uncertain correction factor in any attempt to estimate source densities). Relaxing the requirement to have reliable and physically meaningful estimates of $n_{\rm obs}$, allows us to expand the sample to include additional bursts, FRBs: 20180924, 20190711, 20190110C, 20190425A, 20190520B, 20190303A, 20191106A, 20200223B with well constrained rates and distances (James et al., 2020; Kumar et al., 2021; Niu et al., 2022; Uno et al., 2025; Chawla et al., 2025; Shannon et al., 2025). In the bottom panel of Fig.1 we plot the rate per source ($\mathcal{R}$) directly as a function of their distance ($d_{\rm p}$) without attempting to infer densities from non-nearest members or from surveys with unreported sky coverage. Consistency with the PL distribution of activity rates, $n_{\rm obs}\propto\mathcal{R}^{-a}$ is equivalent to an upper limit in the $d_{\rm p}$-$\mathcal{R}$ plane in which sources are found, which scales as $\mathcal{R}\propto d_{\rm p}^{3/a}$ (by construction, individual sources can be further away from the nearest member of their source sub-class, but not vice versa).
Appendix C Correlation between rate and maximal energy